a2 + b2 + c2 - 2a - 6b - 8c + 26 = 0
find a, b, c

Grade:9

1 Answers

A M S ARUN KRISHNA
8 years ago
a2 + b2 + c2 – 2a – 6b – 8c + 26 = 0
a2 + b2 + c2 – 2a – 6b – 8c + 26 = 0
a2 – 2a + b2 – 6b + c2  – 8c + 26 = 0
(a – 1)2 + (b – 3)2 + (c – 4)2 + 26 = 0
(a – 1)2 = 0
(a – 1) =0
a = 1
like that b = 3, c = 4
 
note:
(a – 1)2 = a2 + 1 + 2a
this 1 is included in that 26
like that only for b and c